OpenStudy (vheah):

Question about derivative test : I'm doing homework where it asks me to find the max or min through a second derivative test. The second derivative of this function however is a positive constant. Does that mean that I have a minimum ?

OpenStudy (faiqraees):

Yep

OpenStudy (vheah):

@FaiqRaees Okay so the problem I have is that there's a 100cm wire is cut into two pieces, one is bent to form the shape of an equilateral triangle and the other a square. if my second derivative is a constant how do I find the where we could cut the wire for the minimum?

OpenStudy (kanwal32):

see u have to put the first derivative test to zero

OpenStudy (faiqraees):

equate dy/dx to 0

OpenStudy (vheah):

Sorry for the typing. I'm on a spazz lol

OpenStudy (vheah):

I already found the first derivative, and when I did the second I got a constant @kanwal32

OpenStudy (kanwal32):

did u got the constant greater than zero or less than zero????

OpenStudy (vheah):

greater than 0

OpenStudy (kanwal32):

then it will give minimum no need to put the values

OpenStudy (vheah):

also got a crit value of 56.5 btw

OpenStudy (vheah):

okay so the minimum on where to cut the wire is what the 2nd derivative is ?

OpenStudy (kanwal32):

no it is the first derivative

OpenStudy (kanwal32):

u have to minimize the area

OpenStudy (kanwal32):

now put dy/dx=0

OpenStudy (vheah):

did u get -100 / sqrt(3) ?

OpenStudy (kanwal32):

sorry i got this wrong

OpenStudy (kanwal32):

i am now doing it correctly

OpenStudy (vheah):

It's okay

OpenStudy (kanwal32):

let side of triangle is x and side of square be y

OpenStudy (kanwal32):

perimeter of triangle + perimeter of square=total wire length

OpenStudy (kanwal32):

are u getting this

OpenStudy (vheah):

okay so like this ? f := ((100-x)*(1/4))^2+(1/4)*sqrt(3)*((1/3)*x)^2

OpenStudy (kanwal32):

4y+3x=100 y=(100-3x)/4

OpenStudy (vheah):

yeah

OpenStudy (kanwal32):

area of square=y^2 (100-3x/4)^2

OpenStudy (vheah):

I got this one for the total length wire :

OpenStudy (kanwal32):

yes u are right

OpenStudy (vheah):

when I set that to 0 I got

OpenStudy (vheah):

So that is where the wire is cut?

OpenStudy (kanwal32):

yes

OpenStudy (vheah):

Awesome. And how di find the max area ?

OpenStudy (kanwal32):

i think your x value is wrong

OpenStudy (vheah):

why? shouldi get more than one answer?

OpenStudy (kanwal32):

can u show me the equation where u put x=0

OpenStudy (vheah):

im doing this through maple so its gnna look kinda weird

OpenStudy (kanwal32):

in the first derivative

OpenStudy (vheah):

I think I figured it out. Thanks for the help @kanwal32 :)

OpenStudy (kanwal32):

if u got x=that then plug into the y equation to get maximum area

OpenStudy (kanwal32):

into the first equaiton

OpenStudy (kanwal32):

@vheah